# "You Are Rejected!": An Empirical Study of Large Language Models Taking Hiring Evaluations

### v1

@@ 4.1 Quantitative Analysis (RQ1)
We conduct a quantitative evaluation of various state-of-the-art (SOTA) LLMs. As summarized in Table 1, all models exhibit a statistically significant discrepancy from human preferences. Specifically, the RMSE of non-reasoning models exceeds 2.0 on a 1-9 Likert scale, reflecting a considerable misalignment between LLM-generated responses and human expectations. Among these models, Gemini2.5-flash achieves the lowest RMSE of $2.1060$. For reasoning models, we evaluate DeepSeek-R1 and Gemini2.5-pro. While Gemini2.5-pro performs similarly to Gemini2.5-flash, we observe that DeepSeek-R1 achieves a markedly lower RMSE of $1.5870$, demonstrating its stronger comprehension of the questionnaire.
@@
The hiring evaluation questions can be categorized into three groups: (1) positive questions, which expect answers in a positive direction; (2) negative questions, which aim at receiving disagreement; (3) self-descriptive questions, which allow candidates to respond based on their self-perception. As shown in Figure 2, we compare the average scores of LLMs and human references on positive and negative questions. While DeepSeek-R1 demonstrates an ability to disagree with negative questions, with an average score below $5$ (Neutral), other models fail to make this distinction apparent. The results indicate that LLMs exhibit a significant tendency to select positive options, highlighting a divergence from human preferences. As illustrated in the Appendix C.1, our findings remain robust even when the option scales are reversed (with 1 for agreement and 9 for disagreement.
@@ C.1 RQ1 Supplement Results
Models P P-revert N N-revert
@@
Table 3: A comparison of indicators with original versus reversed scoring criteria. P represents the average score of all answers to positive questions, and N represents for negative questions. ”revert” indicates the result after the indicator is reversed.
The operation of the indicator revert involves inverting the “1–9 scoring scale and its one-to-one correspondence from strongly disagree to very strongly agree” as defined in the Answer Strategy Instructions part of the prompt (Figure 5). For DeepSeek3.1-Terminus and DeepSeek3.2-Exp, we obtain the results that $S+S_{revert}\in[10,11]$ (S represents P or N) ,demonstrating that the generated responses remain stable even when the evaluation metric is reverted. This consistency effectively rules out the hypothesis that LLM exhibits a systematic preference toward assigning "Positive" scores across the questions.
@@ C.4 Case Study
Combining Table 4 with Figures 7 and 8, we can verify the previous conclusions.

### v2

@@ 4.1 Quantitative Analysis (RQ1)
We conduct a quantitative evaluation of various state-of-the-art (SOTA) LLMs. As summarized in Table 1, all models exhibit a statistically significant discrepancy from human preferences. Specifically, the RMSE of non-reasoning models exceeds $2.0$ on a 11-99 Likert scale, reflecting a considerable misalignment between LLM-generated responses and human expectations. Among these models, Gemini2.5-flash achieves the lowest RMSE of $2.1060$. For reasoning models, we evaluate DeepSeek-R1 and Gemini2.5-pro. While Gemini2.5-pro performs similarly to Gemini2.5-flash, we observe that DeepSeek-R1 achieves a markedly lower RMSE of $1.5870$, demonstrating its stronger comprehension of the questionnaire.
@@
The hiring evaluation questions can be categorized into three groups: (1) positive questions, which expect answers in a positive direction; (2) negative questions, which aim at receiving disagreement; (3) self-descriptive questions, which allow candidates to respond based on their self-perception. As shown in Figure 2, we compare the average scores of LLMs and human references on positive and negative questions. While DeepSeek-R1 demonstrates an ability to disagree with negative questions, with an average score below $5$ (Neutral), other models fail to make this distinction apparent. The results indicate that LLMs exhibit a significant tendency to select positive options, highlighting a divergence from human preferences. As illustrated in the Appendix C.1, our findings remain robust even when the option scales are reversed (with $1$ for agreement and $9$ for disagreement).
@@ C.1 RQ1 Supplement Results
Models P P-reverse N N-reverse
@@
Table 3: A comparison of indicators with original versus reversed scoring criteria. P represents the average score of all answers to positive questions, and N represents for negative questions. "reverse" indicates the result after the indicator is reversed.
The operation of the indicator reverse involves inverting the "1–9 scoring scale and its one-to-one correspondence from strongly disagree to very strongly agree" as defined in the Answer Strategy Instructions part of the prompt (Figure 5). For DeepSeek3.1-Terminus and DeepSeek3.2-Exp, we obtain the results that $S+S_{reverse}\in[10,11]$ ($S$ represents P or N) , demonstrating that the generated responses remain stable even when the evaluation metric is reverseed. This consistency effectively rules out the hypothesis that LLM exhibits a systematic preference toward assigning "Positive" scores across the questions.
@@ C.4 Case Study
Combining Table 4 with the 2 figures mentioned above, we can verify the previous conclusions.